# From Empirical Observations to Universality: Dynamics of Deep Learning with Inputs Built on Gaussian mixture

### v1

@@ B Derivation of the ODE for Covariance and Weights
$\left[\sum_{j=1}^{K}v_{j}g(\lambda_{j})-\sum_{m=1}^{M}\tilde{v}_{m}\tilde{g}(v_{m})\right]g^{\prime}(\lambda_{k})f(U_{i})$ (S30)
@@
$\begin{split}S_{k,r}:=S_{k,r}-\cfrac{\eta}{\sqrt{N}}v^{k}\biggl{[}\sum_{j\neq k}^{K}v_{j}\mathbb{E}[g(\lambda_{j})g^{\prime}(\lambda_{k})\beta_{r}]+v_{k}\mathbb{E}[g(\lambda_{k})g^{\prime}(\lambda_{k})\beta_{r}]\\ -\sum_{n}^{M}\tilde{v}_{n}\mathbb{E}[\tilde{g}(\nu_{n})g^{\prime}(\lambda_{k})\beta_{r}]\biggr{]}\end{split}$ (S31)
@@
$\mathcal{S}_{k,r}=\cfrac{1}{\sqrt{D}}\sum_{r}S_{k,r}(\psi_{\tau})_{r},\quad\mathcal{W}_{m,\tau}=\cfrac{1}{\sqrt{D}}\sum_{r}\tilde{W}_{m,r}(\psi_{\tau})_{r}$ (S33)
@@
$\begin{split}\mathcal{S}_{k,\tau}:=\mathcal{S}_{k,\tau}-\cfrac{\eta}{\sqrt{DN}}v^{k}\sum_{r}(\psi_{\tau})_{r}\biggl{[}\sum_{j\neq k}^{K}v_{j}\mathbb{E}[g(\lambda_{j})g^{\prime}(\lambda_{k})\beta_{r}]+v_{k}\mathbb{E}[g(\lambda_{k})g^{\prime}(\lambda_{k})\beta_{r}]\\ -\sum_{n}^{M}\tilde{v}_{n}\mathbb{E}[\tilde{g}(\nu_{n})g^{\prime}(\lambda_{k})\beta_{r}]\biggr{]}\end{split}$ (S35)
@@
$(R_{k,m})_{t+1}-(R_{k,m})_{t}=\cfrac{b}{D}\sum_{\tau}\left[(\mathcal{S}_{k,\tau})_{t+1}-(\mathcal{S}_{k,\tau})_{t}\right]\tilde{W}_{m,r}$ (S36)
@@
$\mathcal{T}_{m,n}\equiv\cfrac{1}{D}\sum_{\tau}\rho_{\tau}\tilde{W}_{m,r}\tilde{W}_{n,r},\quad\cfrac{1}{D}\sum_{\tau}\rho_{\tau}\mathcal{S}_{\ell,\tau}\tilde{W}_{n,\tau}$ (S37)
@@
$r_{k,m}(\rho)=\cfrac{1}{\epsilon_{\rho}}\cfrac{1}{D}\sum_{\tau}\tilde{S}_{k,\tau}\tilde{W}_{m,\tau}\mathbf{1}_{\rho_{\tau}\in[\rho,\rho+\epsilon_{\rho}]}$ (S38)
@@
$r_{km}(\rho)\sum_{n}^{M}\tilde{v}_{n}\frac{T_{nn}\mathbb{E}[g^{\prime}(\lambda_{k})\lambda_{k}\tilde{g}(\nu_{n})]-R_{kn}\mathbb{E}[g^{\prime}(\lambda_{k})\nu_{n}\tilde{g}(\nu_{n})]}{Q_{kk}T_{nn}-\left(R_{kn}\right)^{2}}$
$-\frac{b\rho}{d(\rho)}\sum_{n}^{M}\tilde{v}_{n}\mathcal{T}_{nm}\frac{Q_{kk}\mathbb{E}[g^{\prime}(\lambda_{k})\nu_{n}\tilde{g}(\nu_{n})]-R_{kn}\mathbb{E}[g^{\prime}(\lambda_{k})\lambda_{k}\tilde{g}(\nu_{n})]}{Q_{kk}T_{nn}-\left(R_{kn}\right)^{2}}\Biggr{)}$
@@
Following a similar process as before, we find that the first term, $W_{k,l}$, adheres to:
$\frac{\mathrm{d}W_{k,\ell}(t)}{\mathrm{d}t}=$ $-\eta v_{k}\left(\sum_{j}^{K}v_{j}\mathbb{E}[g^{\prime}(\lambda_{k})\lambda_{\ell}{g}(\lambda_{j})]-\sum_{n}\tilde{v}_{n}\mathbb{E}[g^{\prime}(\lambda_{k})\lambda_{\ell}\tilde{g}(\nu_{n})]\right)$ (S43)
$-\eta v_{\ell}\left(\sum_{j}^{K}v_{j}\mathbb{E}[g^{\prime}(\lambda_{\ell})\lambda_{k}{g}(\lambda_{j})]-\sum_{n}\tilde{v}_{n}\mathbb{E}[g^{\prime}(\lambda_{\ell})\lambda_{k}\tilde{g}(\nu_{n})]\right)$
@@
$-2\sum_{j}^{K}\sum_{m}^{M}v_{j}\tilde{v}_{m}\mathbb{E}[g^{\prime}(\lambda_{k})g^{\prime}(\lambda_{\ell})g(\lambda_{j})\tilde{g}(\nu_{m})]$
$+\sum_{n,m}^{M}\tilde{v}_{n}\tilde{v}_{m}\mathbb{E}[g^{\prime}(\lambda_{k})g^{\prime}(\lambda_{\ell})\tilde{g}(\nu_{n})\tilde{g}(\nu_{m})]\biggr{)}$
@@
$-d(\rho)v_{k}\sigma_{k\ell}(\rho)\sum_{n}\tilde{v}_{n}\frac{T_{nn}\mathbb{E}[g^{\prime}(\lambda_{k})\lambda_{k}\tilde{g}(\nu_{n})]-R_{kn}\mathbb{E}[g^{\prime}(\lambda_{k})\nu_{n}\tilde{g}(\nu_{n})]}{Q_{kk}T_{nn}-\left(R_{kn}\right)^{2}}$
$-b\rho v_{k}\sum_{n}\tilde{v}_{n}r_{\ell n}(\rho)\frac{Q_{kk}\mathbb{E}[g^{\prime}(\lambda_{k})\nu_{n}\tilde{g}(\nu_{n})]-R_{kn}\mathbb{E}[g^{\prime}(\lambda_{k})\lambda_{k}\tilde{g}(\nu_{n})]}{Q_{kk}T_{nn}-\left(R_{kn}\right)^{2}}$
@@
$\left.-2\sum_{j}^{K}\sum_{m}^{M}v_{j}\tilde{v}_{m}\mathbb{E}[g^{\prime}(\lambda_{k})g^{\prime}(\lambda_{\ell})g(\lambda_{j})\tilde{g}(\nu_{m})+\sum_{n,m}^{M}\tilde{v}_{n}\tilde{v}_{m}\mathbb{E}[g^{\prime}(\lambda_{k})g^{\prime}(\lambda_{\ell})\tilde{g}(\nu_{n})\tilde{g}(\nu_{m})]\right)$
@@
$\frac{dv_{k}}{dt}=\eta\left[\sum^{M}_{n}\tilde{v}_{n}\mathbb{E}[g(\lambda_{k})\tilde{g}(\nu_{n})]-\sum^{K}_{j}v_{j}\mathbb{E}[g(\lambda_{k}){g}(\lambda_{j})]\right]$ (S47)
@@
$\epsilon_{g}(\theta,\widetilde{\theta})$ $=\frac{1}{2}\mathbb{E}\left[\left(\sum_{k}^{K}v_{k}g\left(\lambda_{k}\right)-\sum_{m}^{M}\tilde{v}_{m}\tilde{g}\left(\nu_{m}\right)\right)^{2}\right]$ (S48)
$=\frac{1}{2}\sum_{k,\ell}v_{k}v_{\ell}\mathbb{E}[g(\lambda_{k}){g}(\lambda_{\ell})]+\frac{1}{2}\sum_{n,m}\tilde{v}^{n}\tilde{v}^{m}\mathbb{E}[\tilde{g}(\nu_{n})\tilde{g}(\nu_{m})]-\sum_{k,n}v_{k}\tilde{v}_{n}\mathbb{E}[g(\lambda_{k})\tilde{g}(\nu_{n})]$
C Derivation of Function Expectation Approximation Lemma
@@
The distribution of the sample mean follows a Gaussian distribution under the limit of a large number of samples, $\bar{r}\sim\mathcal{N}(\mu,\sigma)$. Under this, we can empirically verify the lemma results for erf, ReLU, and sgn. The Figure S1 shows approximately the same expectation of sample mean $\mathbb{E}[\overline{f(x)}]\to\mu$.
D KL divergence under Gaussian Mixture setting
Experimentally, by adjusting the distribution bound $\alpha$ of the mixture component means $\mu_{i}$, we observed a regular correlation in the KL divergence between the mixture distribution and the normal Gaussian distribution [Figure S1].
Figure S2: Numerical results depicting the relationship between KL divergence and the parameter $\alpha$ for different numbers of Gaussian mixture components, $m$.
This approach enabled us to conduct empirical analyses on various Gaussian mixtures by adjusting the values of $\alpha$ and $m$.
E Additional Detailed Experimental Conditions
In this study, the $C$ dimension was set to $D=500$ and the $X$ dimension for student input to $N=1000$. The dimensions of the hidden layers for both teacher and student models were uniformly set to $K=M=2$. For the activation functions, both the teacher and student models utilized the same function, $g(x)=\tilde{g}(x)=\text{erf}(x/\sqrt{2})$ or ReLU function. The nonlinear function $f(x)=\text{sgn}(x)$ was employed to generate the student input. The learning rate was set at $\gamma=0.2$, and training was conducted using the Mean Squared Error (MSE) loss, $\mathcal{L}=1/2(\hat{y}-y)^{2}$. The SGD update used was a single batch SGD with layerwise learning rate scaling, as mentioned earlier II.2. For training, the neural network was updated for a total of $100\times 1000$ steps.
For more detailed information on the numerical SGD implementation and ODE update implementation, please refer to the following code repository: https://github.com/peardragon/_GaussianMixture. Our research was carried out on a computing setup equipped with an AMD Ryzen 7 7700X CPU and an NVIDIA GeForce RTX 3060 12GB GPU.
F Detailed tLog-RMSE Results
In the main text, we utilized the average tLog-RMSE for the sake of conciseness and clarity in presenting our results. However, to provide additional detail, this section provides the average values of each dynamics metric and their corresponding tLog-RMSE values. The definitions inherent to each dynamics metric suggest that for $Q$, off-diagonal terms are expected to be less correlated, whereas for $R$, it is the diagonal terms that are anticipated to exhibit less correlation. This expectation is shown in our results, where metrics indicative of less correlation elements demonstrated lower values.
The observations shared in the main body of the text, the number of components $m$ impacts are negligible under the standardization setting. Herein, we focus specifically on the case where $m=2$, presenting the outcomes in a tabular format. Each value has been rounded to three decimal places .
Table 1: $m=2$, erf case
$Q_{0,0}$ $Q_{0,1}$ $Q_{1,0}$ $Q_{1,1}$ $R_{0,0}$ $R_{0,1}$ $R_{1,0}$ $R_{1,1}$ $v_{0}$ $v_{1}$
$\mathcal{M}_{\alpha=0}$ 0.090 0.040 0.035 0.093 0.030 0.036 0.036 0.032 0.012 0.014
$\mathcal{M}_{\alpha=1}$ 0.083 0.044 0.039 0.092 0.028 0.036 0.036 0.032 0.012 0.013
$\mathcal{M}_{\alpha=2}$ 0.091 0.044 0.037 0.092 0.028 0.036 0.036 0.032 0.013 0.015
$\mathcal{M}_{\alpha=3}$ 0.086 0.044 0.039 0.095 0.028 0.035 0.037 0.032 0.012 0.014
$\mathcal{M}_{\alpha=4}$ 0.091 0.041 0.037 0.094 0.028 0.038 0.037 0.033 0.013 0.014
$\mathcal{M}_{\alpha=5}$ 0.085 0.042 0.037 0.088 0.028 0.037 0.036 0.031 0.013 0.014
$\mathcal{M}_{\alpha=6}$ 0.083 0.043 0.036 0.096 0.027 0.034 0.036 0.032 0.012 0.014
$\mathcal{M}_{\alpha=7}$ 0.088 0.044 0.040 0.092 0.027 0.036 0.037 0.032 0.013 0.014
$\mathcal{M}_{\alpha=8}$ 0.090 0.042 0.037 0.094 0.029 0.036 0.037 0.032 0.012 0.014
$\mathcal{M}_{\alpha=9}$ 0.083 0.041 0.037 0.089 0.029 0.035 0.036 0.030 0.013 0.015
$\mathbb{E}[|X|]$ 0.995 0.093 0.093 0.988 0.088 0.126 0.129 0.085 0.488 0.508
Table 2: $m=2$, ReLU case
$Q_{0,0}$ $Q_{0,1}$ $Q_{1,0}$ $Q_{1,1}$ $R_{0,0}$ $R_{0,1}$ $R_{1,0}$ $R_{1,1}$ $v_{0}$ $v_{1}$
$\mathcal{M}_{\alpha=0}$ 0.037 0.036 0.038 0.041 0.024 0.024 0.026 0.028 0.006 0.006
$\mathcal{M}_{\alpha=1}$ 0.037 0.035 0.038 0.041 0.024 0.025 0.027 0.027 0.007 0.006
$\mathcal{M}_{\alpha=2}$ 0.036 0.036 0.038 0.039 0.025 0.025 0.026 0.028 0.006 0.006
$\mathcal{M}_{\alpha=3}$ 0.035 0.035 0.039 0.040 0.025 0.025 0.027 0.028 0.008 0.006
$\mathcal{M}_{\alpha=4}$ 0.037 0.036 0.037 0.044 0.025 0.025 0.026 0.031 0.006 0.006
$\mathcal{M}_{\alpha=5}$ 0.037 0.037 0.037 0.041 0.025 0.026 0.027 0.028 0.006 0.007
$\mathcal{M}_{\alpha=6}$ 0.035 0.036 0.037 0.042 0.025 0.024 0.027 0.028 0.006 0.006
$\mathcal{M}_{\alpha=7}$ 0.037 0.035 0.038 0.041 0.025 0.026 0.027 0.028 0.007 0.007
$\mathcal{M}_{\alpha=8}$ 0.037 0.036 0.038 0.039 0.025 0.024 0.026 0.027 0.006 0.006
$\mathcal{M}_{\alpha=9}$ 0.036 0.033 0.040 0.038 0.026 0.025 0.026 0.028 0.006 0.006
$\mathbb{E}[|X|]$ 0.847 0.053 0.064 1.009 0.053 0.089 0.072 0.165 0.507 0.597
G Results under erf activation function
Similar results can be observed even when using the error function (erf) as the activation function.
Figure S3: Log Sampled MSE under KL divergence in un-standardized Gaussian mixture setting, under erf activation function. a, b, and c display the results corresponding to metrics $Q$, $R$, and $v$, respectively. The shaded areas around each curve indicate the standard error, across varying numbers of Gaussian mixture components $m$.
Figure S4: Temporal log sampled RMSE(tLog-RMSE) under KL divergence in standardized Gaussian mixture setting, under erf activation function. a, b, and c display the results corresponding to metrics $Q$, $R$, and $v$, respectively, with the standard error represented by the shaded regions around each curve, across different numbers of Gaussian mixture components $m$. To illustrate the minimal scale of the error, a dotted line representing the mean of the each dynamic metric values $\mathbb{E}[Q]$, $\mathbb{E}[R]$, $\mathbb{E}[v]$ is added, highlighting the small error magnitude in comparison to the overall scale of the values. Additionally, a randomly sampled example of evolution of dynamic metrics, demonstrating almost negligible error, is included as an inset.

### v2

@@ 3 Derivation of the ODE for Covariance and Weights
$\left[\sum_{j=1}^{K}v_{j}g(\lambda_{j})-\sum_{m=1}^{M}\widetilde{v}_{m}\widetilde{g}(v_{m})\right]g^{\prime}(\lambda_{k})f(U_{i})$ (S30)
@@
$\begin{split}S_{k,r}:=S_{k,r}-\cfrac{\eta}{\sqrt{N}}v^{k}\biggl{[}\sum_{j\neq k}^{K}v_{j}\mathbb{E}[g(\lambda_{j})g^{\prime}(\lambda_{k})\beta_{r}]+v_{k}\mathbb{E}[g(\lambda_{k})g^{\prime}(\lambda_{k})\beta_{r}]\\ -\sum_{n}^{M}\widetilde{v}_{n}\mathbb{E}[\widetilde{g}(\nu_{n})g^{\prime}(\lambda_{k})\beta_{r}]\biggr{]}\end{split}$ (S31)
@@
$\mathcal{S}_{k,r}=\cfrac{1}{\sqrt{D}}\sum_{r}S_{k,r}(\psi_{\tau})_{r},\quad\mathcal{W}_{m,\tau}=\cfrac{1}{\sqrt{D}}\sum_{r}\widetilde{W}_{m,r}(\psi_{\tau})_{r}$ (S33)
@@
$\begin{split}\mathcal{S}_{k,\tau}:=\mathcal{S}_{k,\tau}-\cfrac{\eta}{\sqrt{DN}}v^{k}\sum_{r}(\psi_{\tau})_{r}\biggl{[}\sum_{j\neq k}^{K}v_{j}\mathbb{E}[g(\lambda_{j})g^{\prime}(\lambda_{k})\beta_{r}]+v_{k}\mathbb{E}[g(\lambda_{k})g^{\prime}(\lambda_{k})\beta_{r}]\\ -\sum_{n}^{M}\widetilde{v}_{n}\mathbb{E}[\widetilde{g}(\nu_{n})g^{\prime}(\lambda_{k})\beta_{r}]\biggr{]}\end{split}$ (S35)
@@
$(R_{k,m})_{t+1}-(R_{k,m})_{t}=\cfrac{b}{D}\sum_{\tau}\left[(\mathcal{S}_{k,\tau})_{t+1}-(\mathcal{S}_{k,\tau})_{t}\right]\widetilde{W}_{m,r}$ (S36)
@@
$\mathcal{T}_{m,n}\equiv\cfrac{1}{D}\sum_{\tau}\rho_{\tau}\widetilde{W}_{m,r}\widetilde{W}_{n,r},\quad\cfrac{1}{D}\sum_{\tau}\rho_{\tau}\mathcal{S}_{\ell,\tau}\widetilde{W}_{n,\tau}$ (S37)
@@
$r_{k,m}(\rho)=\cfrac{1}{\epsilon_{\rho}}\cfrac{1}{D}\sum_{\tau}\widetilde{S}_{k,\tau}\widetilde{W}_{m,\tau}\mathbf{1}_{\rho_{\tau}\in[\rho,\rho+\epsilon_{\rho}]}$ (S38)
@@
$r_{km}(\rho)\sum_{n}^{M}\widetilde{v}_{n}\frac{T_{nn}\mathbb{E}[g^{\prime}(\lambda_{k})\lambda_{k}\widetilde{g}(\nu_{n})]-R_{kn}\mathbb{E}[g^{\prime}(\lambda_{k})\nu_{n}\widetilde{g}(\nu_{n})]}{Q_{kk}T_{nn}-\left(R_{kn}\right)^{2}}$
$-\frac{b\rho}{d(\rho)}\sum_{n}^{M}\widetilde{v}_{n}\mathcal{T}_{nm}\frac{Q_{kk}\mathbb{E}[g^{\prime}(\lambda_{k})\nu_{n}\widetilde{g}(\nu_{n})]-R_{kn}\mathbb{E}[g^{\prime}(\lambda_{k})\lambda_{k}\widetilde{g}(\nu_{n})]}{Q_{kk}T_{nn}-\left(R_{kn}\right)^{2}}\Biggr{)}$
@@
Notice that we ignore $a$ term since we focus on symmetric nonlinear function $f$. Following a similar process as before, we find that the first term, $W_{k,l}$, adheres to:
$\frac{\mathrm{d}W_{k,\ell}(t)}{\mathrm{d}t}=$ $-\eta v_{k}\left(\sum_{j}^{K}v_{j}\mathbb{E}[g^{\prime}(\lambda_{k})\lambda_{\ell}{g}(\lambda_{j})]-\sum_{n}\widetilde{v}_{n}\mathbb{E}[g^{\prime}(\lambda_{k})\lambda_{\ell}\widetilde{g}(\nu_{n})]\right)$ (S43)
$-\eta v_{\ell}\left(\sum_{j}^{K}v_{j}\mathbb{E}[g^{\prime}(\lambda_{\ell})\lambda_{k}{g}(\lambda_{j})]-\sum_{n}\widetilde{v}_{n}\mathbb{E}[g^{\prime}(\lambda_{\ell})\lambda_{k}\widetilde{g}(\nu_{n})]\right)$
@@
$-2\sum_{j}^{K}\sum_{m}^{M}v_{j}\widetilde{v}_{m}\mathbb{E}[g^{\prime}(\lambda_{k})g^{\prime}(\lambda_{\ell})g(\lambda_{j})\widetilde{g}(\nu_{m})]$
$+\sum_{n,m}^{M}\widetilde{v}_{n}\widetilde{v}_{m}\mathbb{E}[g^{\prime}(\lambda_{k})g^{\prime}(\lambda_{\ell})\widetilde{g}(\nu_{n})\widetilde{g}(\nu_{m})]\biggr{)}$
@@
$-d(\rho)v_{k}\sigma_{k\ell}(\rho)\sum_{n}\widetilde{v}_{n}\frac{T_{nn}\mathbb{E}[g^{\prime}(\lambda_{k})\lambda_{k}\widetilde{g}(\nu_{n})]-R_{kn}\mathbb{E}[g^{\prime}(\lambda_{k})\nu_{n}\widetilde{g}(\nu_{n})]}{Q_{kk}T_{nn}-\left(R_{kn}\right)^{2}}$
$-b\rho v_{k}\sum_{n}\widetilde{v}_{n}r_{\ell n}(\rho)\frac{Q_{kk}\mathbb{E}[g^{\prime}(\lambda_{k})\nu_{n}\widetilde{g}(\nu_{n})]-R_{kn}\mathbb{E}[g^{\prime}(\lambda_{k})\lambda_{k}\widetilde{g}(\nu_{n})]}{Q_{kk}T_{nn}-\left(R_{kn}\right)^{2}}$
@@
$\left.-2\sum_{j}^{K}\sum_{m}^{M}v_{j}\widetilde{v}_{m}\mathbb{E}[g^{\prime}(\lambda_{k})g^{\prime}(\lambda_{\ell})g(\lambda_{j})\widetilde{g}(\nu_{m})+\sum_{n,m}^{M}\widetilde{v}_{n}\widetilde{v}_{m}\mathbb{E}[g^{\prime}(\lambda_{k})g^{\prime}(\lambda_{\ell})\widetilde{g}(\nu_{n})\widetilde{g}(\nu_{m})]\right)$
@@
$\frac{dv_{k}}{dt}=\eta\left[\sum^{M}_{n}\widetilde{v}_{n}\mathbb{E}[g(\lambda_{k})\widetilde{g}(\nu_{n})]-\sum^{K}_{j}v_{j}\mathbb{E}[g(\lambda_{k}){g}(\lambda_{j})]\right]$ (S47)
@@
$\epsilon_{g}(\theta,\widetilde{\theta})$ $=\frac{1}{2}\mathbb{E}\left[\left(\sum_{k}^{K}v_{k}g\left(\lambda_{k}\right)-\sum_{m}^{M}\widetilde{v}_{m}\widetilde{g}\left(\nu_{m}\right)\right)^{2}\right]$ (S48)
$=\frac{1}{2}\sum_{k,\ell}v_{k}v_{\ell}\mathbb{E}[g(\lambda_{k}){g}(\lambda_{\ell})]+\frac{1}{2}\sum_{n,m}\widetilde{v}^{n}\widetilde{v}^{m}\mathbb{E}[\widetilde{g}(\nu_{n})\widetilde{g}(\nu_{m})]-\sum_{k,n}v_{k}\widetilde{v}_{n}\mathbb{E}[g(\lambda_{k})\widetilde{g}(\nu_{n})]$
B Results: Additional results
Similar convergence results can be observed even when using the error function (erf).
Figure S1: Examples of dynamics under standardized Gaussian mixtures with $m=2,m=16$, and $\alpha=0.001,0.01,0.1,1,10$. Dynamics of (a) generalization error $\epsilon_{g}$, (b) covariance matrix $Q$, (c) covariance matrix $R$, and (d) weight of the second layer $v$. The SGD results are averaged over 10 runs. Using error function for teacher and student model’s activation function. Notice that error function’s bound property ($\text{erf}(\cdot)\in(-1,1)$) satisfy lemmma V.3 in loose manner.
C Discussion: Additional derivation and results in discussion
1 Derivation of Function Expectation Approximation Lemma
@@
The distribution of the sample mean follows a Gaussian distribution under the limit of a large number of samples, $\overline{r}\sim\mathcal{N}(\mu,\sigma)$. Under this, we can empirically verify the lemma results for erf, ReLU, and sgn. The Fig. S2 shows approximately the same expectation of sample mean $\mathbb{E}[\overline{f(x)}]\to\mu$.
2 Discussion: Additional Results from various distribution settings
The nonlinear function $f=tanh(\cdot)$ also satisfying condition, yield similar convergence phenomena.
Figure S3: Examples of dynamics under various distribution settings. Dynamics of (a) generalization error $\epsilon_{g}$, (b) covariance matrix $Q$, (c) covariance matrix $R$, and (d) weight of the second layer $v$. The SGD results are averaged over 10 runs. Using $tanh$ function for nonlinear function. Note that even the nonlinear function preserves the characteristics of the input to some extent, the convergence occur.
3 Additional information in pseudo-real dataset
3.1 Experimental Condition
In this pseudo-real dataset investigation, the dimension of teacher model input $C$ was set to $D=500$, and the dimension of student model input $X$ was set to $N=28^{2}$, MNIST dimension. The dimensions of the hidden layers for both teacher and student models were uniformly set to $K=M=2$. Both the teacher and student models employed the same function, $\text{hardtanh}(x)=\max(-1,\min(1,x))$. The nonlinear function $f(x)$ used to generate the student input was $f(\cdot)=\text{hardtanh}(\cdot)$. In $C$ optimization process, we use $10000$ MNIST dataset to find optimal $C$ under loss function. Then, we use kernel density estimation with bandwidth $0.01$ and Gaussian kernel. For teacher model optimization for label matching, we use $10^{6}$ dataset sampled from estimated distribution. The remainder of the process, including SGD and ODE computations, is consistent.
3.2 Results with different teacher model
In this section we present several similar results under different teacher model. All teacher model has accuracy over $60\%$. These accuracy is statistically significant based on a t-test comparison with randomly initialized teacher models.
Figure S4: Examples of dynamics under the pseudo-real MNIST dataset. SGD results are plotted as dashed lines. Both SGD and ODE results are averaged over 10 runs. Trained teacher model accuracy is $61.31\%$.
Figure S5: Examples of dynamics under the pseudo-real MNIST dataset. SGD results are plotted as dashed lines. Both SGD and ODE results are averaged over 10 runs. Trained teacher model accuracy is $61.00\%$.
Figure S6: Examples of dynamics under the pseudo-real MNIST dataset. SGD results are plotted as dashed lines. Both SGD and ODE results are averaged over 10 runs. Trained teacher model accuracy is $67.31\%$.